OpenStudy (anonymous):

an investment is worth $2633 in 1994. by 1998 it has grown to $4061. let y be the value of the investment in the year x, where x=0 represents 1994. write a linear equation that models the value of the investment in the year x. show work!!

OpenStudy (anonymous):

A.) y= 1/357x + 2633 B.) y= -357x + 5489 C.) y= 357x + 2633 D.) y= -357x +2633

OpenStudy (anonymous):

Think it's D, what u think?

OpenStudy (anonymous):

no clue, did you show any work in figuring that out?

OpenStudy (anonymous):

you know point slope formula??????

OpenStudy (anonymous):

NOOOOOOOO it's C for sure. :P

OpenStudy (anonymous):

SLOPE of this points. ie (0,2633) and (4,4061) are